OpenStudy (anonymous):

lim as t approaches 0 of [(1/t(1+t)^1/2) -(1/t)]

OpenStudy (tkhunny):

Have you considered adding the fractions? \(\dfrac{1 - \sqrt{1+t}}{t\cdot\sqrt{1+t}}\) Couple things that could be done, after that. 1) "Rationalize" the Numerator. 2) l'Hospital's Rule for form 0/0.

OpenStudy (anonymous):

Can you please show me how you got to that fraction right there. Just to clarify the original form is \[\frac{ 1 }{ t \sqrt{t+1} }-\frac{ 1 }{ t}\] If that is what you used i'm confused how you got there.

OpenStudy (dumbcow):

he/she combined fractions \[\frac{1}{t \sqrt{1+t}}-\frac{1 \sqrt{1+t}}{t \sqrt{1+t}} = \frac{1-\sqrt{1+t}}{t \sqrt{1+t}}\]

OpenStudy (anonymous):

oh ok you found a common denominator my bad i'm dumb

OpenStudy (tkhunny):

Common Denominator? You've been doing it for a long time. Don't forget things that you should know. Truthfully, it's common for things you know to abandon you when you encounter something new. You have to fight it!

OpenStudy (anonymous):

ok and then next would i multiply by conjugate?

OpenStudy (dumbcow):

yes that would be my suggestion

OpenStudy (anonymous):

ok thanks guys

OpenStudy (dumbcow):

you should end up with limit of -1/2

OpenStudy (anonymous):

awesome thanks